OpenStudy (anonymous):

Would the zeros of the equation 2x^4-5x^3-3x^2=0 be -2,5,-3?

OpenStudy (barrycarter):

What about 0 itself?

OpenStudy (anonymous):

Would that be in with the zeros of the function?

OpenStudy (barrycarter):

What you have is actually an equation, but when x=0, what does the function become? Also, it should have only 2 other 0's, not 3.

OpenStudy (anonymous):

okay so its and equation... I dont quite understand how to do this. It says to solve on calculator and find the zeros... I tried and only got -2,5, and 3.

jimthompson5910 (jim_thompson5910):

how did you get -2 and 5?

OpenStudy (anonymous):

Laughs/... I honestly do not know :/

OpenStudy (barrycarter):

Are you studying algebra or pre-algebra?

OpenStudy (anonymous):

Algebra II

jimthompson5910 (jim_thompson5910):

2x^4-5x^3-3x^2=0 x^2(2x^2 - 5x - 3) = 0 x^2(2x^2 - 6x + x - 3) = 0 x^2(2x(x - 3) + 1(x - 3)) = 0 x^2(2x + 1)(x - 3) = 0 Can you take it from here?

OpenStudy (anonymous):

May you explain what you are doing? Pleease?

jimthompson5910 (jim_thompson5910):

Sure thing In step 2, I'm factoring out the GCF x^2 Then I broke up -5x into -6x + x. This is because -6 and 1 multiply to -6 and add to -5 (notice how 2 times -3 = -6) From there, I'm factoring by grouping

jimthompson5910 (jim_thompson5910):

Does that clear things up?

OpenStudy (anonymous):

Oh okay i get that now...

jimthompson5910 (jim_thompson5910):

alright, tell me what you get for the answer

OpenStudy (anonymous):

So after that last part What would i do i am so so sorry lol

jimthompson5910 (jim_thompson5910):

it's ok

jimthompson5910 (jim_thompson5910):

use the zero product property x^2(2x + 1)(x - 3) = 0 x^2 = 0 or 2x + 1 = 0 or x - 3 = 0 x = 0 or 2x = -1 or x = 3 x = 0 or x = -1/2 or x = 3 Keep in mind that the x^2 = 0 means that you'll have the root x = 0 with a multiplicity of 2

OpenStudy (anonymous):

I just need explantation i get confused easy

jimthompson5910 (jim_thompson5910):

Let me know if that makes sense or not

OpenStudy (anonymous):

Okay so... the GCF is x^2 and you would factor the whole thing by regrouping... got that so far... then you would find the zeros with, the zero product property okay after find the ones that equal to zero and those are your answers?

jimthompson5910 (jim_thompson5910):

yes the term 'zeros' are a bit confusing, so you can call them roots or solutions so when they say "find the zeros", they mean "find the roots or solutions" The roots/solutions in this case are: x = 0, x = -1/2, or x = 3

OpenStudy (anonymous):

Oh Okay i see i get it now i always wonder about roots vs. zeros.... Understanding now thank you!

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

Your the best... (:

jimthompson5910 (jim_thompson5910):

thanks for thinking so, always glad to help

OpenStudy (anonymous):

Your welcome :) <3